Chhom, Inc., manufactures and sells two products: Product F9 and Product U4. Data concerning the expected production of each product and the expected total direct labor-hours (DLHs) required to produce that output appear below:

Expected Production Direct Labor-Hours Per Unit Total Direct Labor-Hours
Product F9 300 6.0 1,800
Product U4 600 3.0 1,800
Total direct labor-hours 3,600

The direct labor rate is $27.80 per DLH. The direct materials cost per unit is $271.90 for Product F9 and $272.20 for Product U4.

The company is considering adopting an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Estimated Expected Activity
Activity Cost Pools Activity Measures Overhead Cost Product F9 Product U4 Total
Labor-related DLHs $ 38,988 1,800 1,800 3,600
Production orders orders 61,910 400 600 1,000
Order size MHs 126,150 3,800 3,700 7,500
$ 227,048

If the company allocates all of its overhead based on direct labor-hours using its traditional costing method, the overhead assigned to each unit of Product U4 would be closest to: (Round your intermediate calculations to 2 decimal places.)

rev: 03_24_2018_QC_CS-119201

Multiple Choice

  • $185.73 per unit

  • $189.21 per unit

  • $32.49 per unit

  • $50.46 per unit
Answer #1

Predetermined Overhead rate = Estimated overhead / Estimated total labor hours

= 227,048 / 3600 = 63.07

Total Overhead allocated to product U4 = 63.07*1800 = 113,526

Per unit allocation = 113,526/600

= 189.21

Option B :D